### v1

# A Unified Bilevel Model for Adversarial Learning and A Case Study

Abstract
Adversarial learning has been attracting more and more attention thanks to the fast development of machine learning and artificial intelligence. However, due to the complicated structure of most machine learning models, the mechanism of adversarial attacks is not well interpreted. How to measure the effect of attack is still not quite clear. In this paper, we propose a unified bilevel model for adversarial learning. We further investigate the adversarial attack in clustering models and interpret it from data perturbation point of view. We reveal that when the data perturbation is relatively small, the clustering model is robust, whereas if it is relatively large, the clustering result changes, which leads to an attack. To measure the effect of attacks for clustering models, we analyse the well-definedness of the so-called $\delta$-measure, which can be used in the proposed bilevel model for adversarial learning of clustering models.
keywords:
convex clustering; adversarial learning; perturbation analysis; robustness; calmness; bilevel optimization
1 Introduction
With the fast development of machine learning and artificial intelligence, adversarial learning is receiving growing attentions. The first striking demonstration came from Szegedy et al. [1], who showed that imperceptible perturbations can reliably force modern neural networks to misclassify inputs, thereby exposing a fundamental vulnerability of high-capacity models. Goodfellow et al. [2] provided a concise explanation and practical attack algorithm so called the Fast Gradient Sign Method (FGSM) for fast, effective adversarial example synthesis, and thus enabled a set of adversarial-training defenses. Kurakin et al. [3] showed that adversarial training can be implemented on larger-scale datasets such as ImageNet, and further revealed that this approach leads to a significant improvement in the robustness of one-step methods. Roberts and Smyth [4] analyzed stochastic gradient descent under Byzantine adversaries, highlighting robustness issues that also arise in distributed or large-scale settings. Chhabra et al. [5] presented a black-box adversarial attack algorithm for clustering models with linearly separable clusters. Later Chhabra et al. [6] proposed a black-box adversarial attack against deep clustering models . We refer to [7, 8, 9] for the review and monographs of adversarial learning.
Following widely adopted taxonomies, adversarial attacks can be divided into three types based on attacker’s knowledge: white-box (full access to parameters/gradients), black-box (only queries or input–output pairs), and gray-box in between. Due to different goals of attacks, it can also be classified as confidence reduction, untargeted misclassification and targeted misclassification. Madry et al. [10] casted adversarial training as min–max robust optimization and used multi-step PGD as a strong white-box baseline for $l_{p}$-bounded threats. Dong et al. [11] added momentum to iterative gradients to markedly improve black-box transferability over standard iterative attacks. Brendel et al. [12] proposed boundary attack that needs only top-1 decisions, walking along the boundary while shrinking perturbations. Papernot et al. [13] targeted a specific label by perturbing a small, saliency chosen set of pixels (sparse $L_{0}$-style attack). Eykholt et al. [14] showed sticker-based perturbations that reliably fool traffic-sign recognition in the real world.
While most exciting literature focuses on developing efficient algorithms to solve adversarial learning model, understanding the mechanism is also extremely crucial in order to improve the robustness of learning models. Below we briefly review some related work that motivates our work in this paper. Moosavi-Dezfooli et al. [15] introduced the DeepFool attack, viewing adversarial perturbations as minimal crossings of local decision boundaries. By iteratively linearizing the classifier, it linked perturbations to decision region geometry and provided a principled way to approximate the smallest such perturbations. Carlini and Wagner [16] formalized perturbation attacks as optimization problems with tailored loss functions, showing attack strength depends critically on objective function and constraint choices. Ilyas et al. [17] offered a feature-based view, arguing that adversarial perturbations exploit non-robust features: predictive yet human-imperceptible statistical patterns. This shifted the focus from geometry to data representation. Su, Li and Cui [18] systematically studied three types of adversarial perturbations, deriving the explicit solutions for sample-adversarial perturbations (sAP), class-universal adversarial perturbations (cuAP) and universal adversarial perturbations (uAP) for binary classification, and approximate the solution for uAP multi-classification case. Later Su and Li [19] addressed the difficulty of generating sAP for nonlinear SVMs via implicit mapping by transforming the perturbation optimization into a solvable nonlinear KKT system.
On the other hand, clustering is a popular yet basic model in machine learning. Various approaches have been proposed to solve clustering problems, including K-means [20, 21], K-medoids [22, 23], hierarchical clustering [24, 25], convex clustering [26, 27, 28, 29, 30] and so on. Among them, convex clustering model [31] attracts great interest due to several advantages, such as the uniqueness of solution and theoretical guarantee of cluster recovery. To deal with high dimensional data clustering, Yuan et al. [32] proposed a dimension reduction technique for structured sparse optimization problems. Ma et al. [33] proposed an improved robust sparse convex clustering (RSCC) model, which incorporates a novel norm-based feature normalization technique to effectively identify and eliminate outlier features. Angelidakis et al. [34] developed improved algorithms for stable instances of clustering problems with center-based objectives, including K-means, K-median and K-center. However, the adversarial attack case for clustering remains untouched.
To summarize, due to the complicated structure of different learning models, most adversarial attacks are difficult to interpret. Therefore, a natural question is whether we can understand the attack for a simple learning model. This motivates our work. In this paper, we study the mechanism of adversarial attack on clustering models. The contribution of the paper can be summarized in three folds. Firstly, we start with the learning model and address the perturbation of learning models. We measure the changes in solution set of learning models under perturbation in terms of calmness. Taking convex clustering as an example, we show that attack can be viewed as a special way of noise which is relatively large. By a 2-way clustering example in the one-dimensional case, we observe that, for clustering problem with noised data, if the noise is relatively small, the clustering result will remain the same. When the noise is relatively large, the clustering result will change, meaning that attack happens. Secondly, we propose the unified bilevel optimization framework for adversarial learning, under which, the adversarial learning can be viewed as seeking the solutions of bilevel models. Finally, we discuss the properties of the deviation function, which is used to measure the effect of adversarial attack. We show the well-definedness of the so-called $\delta$-measure function, and illustrate by different 2-way and 3-way clustering examples.
The organization of the paper is as follows. In Section 2, we investigate the perturbation of learning models and relate the sensitivity of solution set to the so-called calmness property in the context of perturbation analysis. In Section 3, we study the effect of perturbation on the convex clustering model and provide some examples. In Section 4, we propose the unified bilevel optimization model for adversarial learning. In Section 5, we study the so-called $\delta$-measure function to show the well-definedness as the measure of adversarial attack. Final conclusions are given in Section 6.
Notations. We use $\|\cdot\|$ as $l_{2}$ norm for vectors and Frobenius norm for matrices. We use $|V|$ to denote the number of elements in a set $V$.
2 Perturbation of Learning Models
In this part, we will start with the learning model, based on which we will address the perturbation of learning models.
2.1 Learning Model
Let $X\subseteq\mathcal{X}$ be the training data, $Y\subseteq\mathcal{Y}$ be the model parameter in a learning model. The learning process (also referred to as training process) is to find the model parameter $Y^{*}$ by solving the following model:
$\min_{Y\in F(X)}L(X,Y)$ (P)
where $L(X,Y)$ is the objective of the training model, $F(X)\subseteq\mathcal{Y}$ is the feasible set of $Y$, which may be affected by the training data $X$. Let $Y^{*}$ be the optimal solution of (P), which may not be unique. The solution set of (P) is denoted as $S$.
Let the decision function be $D_{Y^{*}}(\cdot)$, where $D_{Y^{*}}(x)$ gives the final decision of a new data $x$. For example, for binary classification model, $D_{Y^{*}}(x)$ is the sign function which gives the label of data $x$, by applying the learning result $Y^{*}$. Specifically, we give two simple examples below.
Example 1.
Support Vector Machine (SVM) for binary classification [35], where $X=\left[\begin{pmatrix}x_{1}\\ y_{1}\end{pmatrix},\cdots,\begin{pmatrix}x_{n}\\ y_{n}\end{pmatrix}\right]\in\mathbb{R}^{(d+1)\times n},\ Y=(\omega,b)\in\mathbb{R}^{d+1}$, with the learning model
$\min_{(\omega,b)\in\mathbb{R}^{d+1}}\quad\frac{1}{2}\|\omega\|^{2}+C\sum_{i=1}^{n}\left(\max\left(0,1-y_{i}(\omega^{\top}{x}_{i}+b)\right)\right)^{2}:=L^{SVM}(X,Y).$ ($l_{2}$-SVM)
The optimal solution of ($l_{2}$-SVM) is denoted as $(\omega^{*},b^{*})$. The decision function $D^{SVM}_{(\omega^{*},b^{*})}(x)=\text{sign}({\omega^{*}}^{\top}x+b^{*})\in\{-1,1\}$.
Example 2.
Convex Clustering [31], where $X=[x_{1},\cdots,x_{n}]\in\mathbb{R}^{d\times n}$, $Y=[y_{1},\cdots,y_{n}]\in\mathbb{R}^{d\times n}$ with the learning model
$\min\limits_{Y\in\mathbb{R}^{d\times n}}L^{CVC}(X,Y)=\frac{1}{2}\sum\limits_{i=1}^{n}\|y_{i}-x_{i}\|^{2}+\gamma\sum\limits_{1\leq i<j\leq n}w_{ij}\|y_{i}-y_{j}\|_{p}$ (CVC)
where $\gamma>0,\ w_{ij}\geq 0,\ i,j=1,\cdots,n$ are given. The optimal solution of (CVC) is $Y^{*}=[y_{1}^{*},\cdots,y_{n}^{*}]\in\mathbb{R}^{d\times n}$. The decision function (i.e., the clustering result) is given by
$D^{CVC}_{Y^{*}}=\{V_{1},\cdots,V_{K}\},$
where $\{V_{1},\cdots,V_{K}\}$ is a partition of $\{1,\cdots,n\}$ and $x_{i}$ and $x_{j}$ are in the same partition if $y_{i}^{*}=y_{j}^{*}$. Since $L^{CVC}(\cdot,\cdot)$ is also strongly convex in $Y$, $Y^{*}$ is the unique solution of (CVC). That is, $S$ is a single singleton.
Remark. Note that for most complicated learning models such as convolutional neural network (CNN), it is usually difficult to write the explicit form of $L(X,Y)$ as well as the decision function $D_{Y^{*}}$. It then brings the challenge in solving the adversarial learning model and interpreting the role of adversarial attack.
2.2 Perturbation of Learning Models
Having introduced the learning model, we are ready to consider the perturbation of the learning model. For general optimization problems, the perturbation analysis is fully addressed in [36]. In the case of noise, let $X(\varepsilon)$ be the noised training data set where $\varepsilon\in\mathcal{X}$ is the perturbation (or noise), the output of the learning model under the perturbation $\varepsilon$ is as follows
$\min_{Y\in F(X(\varepsilon))}L(X(\varepsilon),Y)$ (Pε)
where $Y^{*}(\varepsilon)$ denotes the optimal solution of the learning model under perturbation $\varepsilon$ and $S(\varepsilon)$ denotes the solution set of (Pε).
It is obvious that $X(0)=X$ and $S(0)=S$. Here we would like to highlight that due to different learning models, the form of $\varepsilon$ could be different. If $\mathcal{X}=\mathbb{R}^{k\times n}$, then $\varepsilon$ could be the additive noise or multiplicative noise. However, if $\mathcal{X}$ is in the graph space in graph clustering, i.e., $\mathcal{X}=\left\{\mathcal{G}=(V,E)\ |\ |V|=n,\ E\text{ is any set of edges on }V\right\}$ , then $X(\varepsilon)$ could be the graph that removing vertices or changing set of edges from the current graph $X$.
Intuitively, if the perturbation $\varepsilon$ on data is relatively small, $Y^{*}(\varepsilon)$ may still be different from $Y^{*}$  [^1] Note that $Y^{*}(0)=Y^{*}$. For simplicity, we always use $Y^{*}$ instead of $Y^{*}(0)$., leading to possibly small changes in $S(\varepsilon)$ compared to $S$. If we take $S(\varepsilon)$ as a multifunction of $\varepsilon$, one way to measure the changes of the solution set $S(\varepsilon)$ under the perturbation $\varepsilon$ is calmness, which is a useful property in perturbation analysis. We give the definition below.
Definition 1 (Calmness).
[37, Definition 2] Let $S(\varepsilon)=\arg\min_{Y\in F(X(\varepsilon))}L(X(\varepsilon),Y)$ be a multifunction with a closed graph, denoted as $\operatorname{gph}S(\varepsilon)$, and $(\bar{\varepsilon},\overline{Y})\in\operatorname{gph}S(\varepsilon)$. We say that $S(\cdot)$ is calm at $(\bar{\varepsilon},\overline{Y})$ provided that there exist neighborhoods $\mathcal{N}_{\varepsilon}$ of $\bar{\varepsilon}$ and $\mathcal{N}_{Y}$ of $\overline{Y}$, and a modulus $\mathcal{L}\geq 0$, such that
$S(\varepsilon)\cap\mathcal{N}_{Y}\subset S(\bar{\varepsilon})+\mathcal{L}\|\varepsilon-\bar{\varepsilon}\|\\ \mathbb{B}\quad\text{for all }\varepsilon\in\mathcal{N}_{\varepsilon},$
where $\mathbb{B}$ denotes the closed unit ball in $\mathcal{X}$ and $\operatorname{gph}S(\varepsilon):=\{(\varepsilon,Y):Y\in S(\varepsilon)\}$.
Based on the definition of calmness, one can see that the calmness of $S(\cdot)$ at $\varepsilon=0$ is particularly useful for measuring the changes of $S(\varepsilon)$ relative to $S(0)$. We formally give it below.
Definition 2 (Calmness at $\varepsilon=0$}).
$S(\cdot)$ is calm at $(0,\overline{Y})$ if there exist neighborhoods $\mathcal{N}_{\varepsilon}$ of 0, $\mathcal{N}_{Y}$ of $\overline{Y}$, and a modulus $\mathcal{L}_{0}>0$ such that
$S(\varepsilon)\cap\mathcal{N}_{Y}\subset S+\mathcal{L}_{0}\|\varepsilon\|\ \mathbb{B}\quad\text{for all }\varepsilon\in\mathcal{N}_{\varepsilon}.$
Therefore, one can see that if $S(\cdot)$ is calm at 0, then the changes in the solution set $S(\varepsilon)$ can be controlled by the changes in $\|\varepsilon\|$ (up to the scalar $\mathcal{L}_{0}$). In other words, the robustness of the learning model $L(\cdot,\cdot)$ is closely related to the calmness of the solution set $S(\cdot)$. For a set $S(\cdot)$, there are various ways to check calmness [38, 39, 40]; moreover, Zhou and So [41] provided an equivalent characterization by showing that the error bound property holds if and only if a suitably defined solution mapping is calm, and we will not discuss the details here.
Having successfully measured the changes of $S(\varepsilon)$ due to the perturbation $\varepsilon$, we move on to see whether there is any change in the decision function $D_{Y^{*}}(\cdot)$. Even when the solution set $S(\varepsilon)$ changes, it is still possible that the decision function remains the same. The reason is as follows. In many learning tasks, the decision function is discontinuous. For example, binary classification as shown in Example 1. In other words, for such situation, the perturbation $\varepsilon$ does not have effect on the learning result. We can regard $\varepsilon$ as a neglectable noise in this case. Therefore, the question we would like to ask is as follows: under what condition on $\varepsilon$, the decision function is not changed, i.e., $D_{Y^{*}(\varepsilon)}=D_{Y^{*}}$? The question is not easy to answer. As we mentioned before, this is also highly related to the specific form of the learning model. We will address this question in Section 3, by looking at the convex clustering model as an example.
3 Perturbation Analysis for Convex Clustering
In this section, we take clustering as an example to study the effect of perturbation. Specifically, due to the strongly convex nature of convex clustering model as well as the theoretical recovery result in [31], we choose the convex clustering model (CVC) in Example 2 as our learning model. We start with the case where the small perturbation will not change the clustering result.
Let $X=[x_{1},\cdots,x_{n}]$ be the data and $\mathcal{V}=\{V_{1},\dots,V_{K}\}$ be a partitioning of $X$ and $K$ is the number of clusters. The index sets are defined by
$I_{\alpha}=\left\{i\mid{x}_{i}\in V_{\alpha}\right\},\ n_{\alpha}=|I_{\alpha}|,\ \text{ for }\alpha=1,2,\ldots,K,$
${x}^{(\alpha)}=\frac{1}{n_{\alpha}}\sum_{i\in I_{\alpha}}{x}_{i},\quad w^{(\alpha,\beta)}=\sum_{i\in I_{\alpha}}\sum_{j\in I_{\beta}}w_{ij},\quad\forall\alpha,\beta=1,\ldots,K,$
$w_{i}^{(\beta)}=\sum_{j\in I_{\beta}}w_{ij},\quad\forall i=1,\ldots,n,\ \beta=1,\ldots,K.$
The following result shows the exact recovery result of the learning model (CVC).
Theorem 1.
[31, Theorem 5] Consider the input data $X=\left[{x}_{1},\cdots,{x}_{n}\right]\in\mathbb{R}^{d\times n}$ and its partitioning $\mathcal{V}=\left\{V_{1},V_{2},\ldots,V_{K}\right\}$. Assume that all centroids $\left\{{x}^{(1)},{x}^{(2)},\ldots,{x}^{(K)}\right\}$ are distinct. Let $q\geq 1$ be the conjugate index of $p$ such that $\frac{1}{p}+\frac{1}{q}=1$. $Y^{*}$ is the unique solution of (CVC) and define the map $f\left({x}_{i}\right)={y}_{i}^{*}$ for $i=1,\cdots,n$. Let
$\mu_{ij}^{(\alpha)}:=\sum_{\beta=1,\ \beta\neq\alpha}^{K}\left|w_{i}^{(\beta)}-w_{j}^{(\beta)}\right|,\quad i,j\in I_{\alpha},\ \alpha=1,2,\ldots,K.$
Assume that
$w_{ij}>0\ \text{and}\ n_{\alpha}w_{ij}>\mu_{ij}^{(\alpha)}\ \text{for\ all}\ i,j\in I_{\alpha},\ \alpha=1,\ldots,K.$ (C1)
Let
$\gamma_{\min}:=\max_{1\leq\alpha\leq K}\max_{i,j\in I_{\alpha}}\left\{\frac{\left\|{x}_{i}-{x}_{j}\right\|_{q}}{n_{\alpha}w_{ij}-\mu_{ij}^{(\alpha)}}\right\},\ \ \gamma_{\max}:=\min_{1\leq\alpha<\beta\leq K}\left\{\frac{\left\|{x}^{(\alpha)}-{x}^{(\beta)}\right\|_{q}}{\frac{1}{n_{\alpha}}\sum\limits_{1\leq l\leq K,l\neq\alpha}w^{(\alpha,l)}+\frac{1}{n_{\beta}}\sum\limits_{1\leq l\leq K,l\neq\beta}w^{(\beta,l)}}\right\}.$
If
$\gamma_{\text{min}}<\gamma_{\text{max}}$ (C2)
and $\gamma$ is chosen such that $\gamma\in\left[\gamma_{\min},\gamma_{\max}\right)$, then the map $f$ perfectly recovers $\mathcal{V}$.
Theorem 1 shows that if conditions (C1) and (C2) hold, then the clustering result $D^{CVC}_{Y^{*}}$ coincides with the grand truth partition of $X$. That is, $D^{CVC}_{Y^{*}}=\mathcal{V}$. One can see that the exact recovery is based on conditions (C1) and (C2). Moreover, (C1) and (C2) are calculated based on the ground truth partition $\mathcal{V}$ as well as the data $X$. Given the perturbed data $X(\varepsilon)$, one can make use of Theorem 1 and provide a sufficient condition under which the clustering result is unchanged under perturbation $\varepsilon$. To that end, let $X(\varepsilon)=[x_{1}(\varepsilon),\cdots,x_{n}(\varepsilon)]$ be the perturbation of $X\in\mathbb{R}^{d\times n}$ with $\varepsilon\in\mathbb{R}^{d\times n}$. Under the partition of $D^{CVC}_{Y^{*}}=\left\{V_{1},V_{2},\ldots,V_{K}\right\}$, we define the following notations
${x}^{(\alpha)}({\varepsilon})=\frac{1}{n_{\alpha}}\sum_{i\in I_{\alpha}}{x}_{i}(\varepsilon),\quad w^{(\alpha,\beta)}({\varepsilon})=\sum_{i\in I_{\alpha}}\sum_{j\in I_{\beta}}w_{ij}({\varepsilon}),\quad\forall\alpha,\beta=1,\ldots,K,$ (1)
$w_{i}^{(\beta)}({\varepsilon})=\sum_{j\in I_{\beta}}w_{ij}({\varepsilon}),\quad\forall i=1,\ldots,n,\ \beta=1,\ldots,K.$
Here we use ${x}^{(\alpha)}({\varepsilon}),\ w^{(\alpha,\beta)}({\varepsilon})$ amd $w_{i}^{(\beta)}({\varepsilon})$ to mean that those coefficients may be related to the perturbation $\varepsilon$.
Theorem 2.
Consider the perturbed data $X(\varepsilon)=[x_{1}(\varepsilon),\dots,x_{n}(\varepsilon)]\in\mathbb{R}^{d\times n}$ and the partitioning $D^{CVC}_{Y^{*}}=\left\{V_{1},V_{2},\ldots,V_{K}\right\}$. Let ${x}^{(\alpha)}({\varepsilon})$, $w^{(\alpha,\beta)}({\varepsilon})$ and $w_{i}^{(\beta)}({\varepsilon})$ be defined as in (1). Assume that all centroids $\left\{x^{(1)}(\varepsilon),\dots,x^{(K)}(\varepsilon)\right\}$ are distinct. Let $q\geq 1$ be the conjugate index of $p$ such that $\frac{1}{p}+\frac{1}{q}=1$. Let $Y^{*}(\varepsilon)=[y^{*}_{1}(\varepsilon),\dots,y^{*}_{n}(\varepsilon)]$ be learned via (CVC) in Example 2 and $f_{\varepsilon}:X(\varepsilon)\to Y^{*}(\varepsilon)$ is given by $f_{\varepsilon}(x_{i}\left(\varepsilon)\right)=y^{*}_{i}(\varepsilon)$. Let $\mu_{ij}^{(\alpha)}({\varepsilon}):=\sum\limits_{\beta=1,\beta\neq\alpha}^{k}\left|w_{i}^{(\beta)}({\varepsilon})-w_{j}^{(\beta)}({\varepsilon})\right|,\ i,j\in I_{\alpha},\ \alpha=1,\dots,K$. Assume that
$w_{ij}(\varepsilon)>0\ \text{and}\ n_{\alpha}w_{ij}(\varepsilon)>u_{ij}^{(\alpha)}(\varepsilon)\ \text{for all}\ i,j\in I_{\alpha},\ \alpha=1,\dots,K.$ (C1)
Let
$\gamma_{\text{min}}^{\varepsilon}:=\max\limits_{1\leq\alpha\leq K}\max\limits_{i,j\in I_{\alpha}}\left\{\frac{\|x_{i}(\varepsilon)-x_{j}(\varepsilon)\|_{q}}{n_{\alpha}w_{ij}(\varepsilon)-\mu_{ij}^{(\alpha)}(\varepsilon)}\right\},$
$\gamma_{\text{max}}^{\varepsilon}:=\min\limits_{1\leq\alpha<\beta\leq K}\left\{\frac{\|{x}^{(\alpha)}(\varepsilon)-{x}^{(\beta)}(\varepsilon)\|_{q}}{\frac{1}{n_{\alpha}}\sum\limits_{1\leq l\leq K,l\neq\alpha}w^{(\alpha,l)}(\varepsilon)+\frac{1}{n_{\beta}}\sum\limits_{1\leq l\leq K,l\neq\beta}w^{(\beta,l)}(\varepsilon)}\right\}.$
If
$\gamma_{\text{min}}^{\varepsilon}<\gamma_{\text{max}}^{\varepsilon}$ (C2)
and $\gamma$ is chosen such that $\gamma\in[\gamma_{\text{min}}^{\varepsilon},\gamma_{\text{max}}^{\varepsilon})$, then $D^{CVC}_{Y^{*}(\varepsilon)}=D^{CVC}_{Y^{*}}$, i.e., the clustering result is unchanged.
Proof. By applying [31, Theorem 5 ] with the ground truth partitioning $\mathcal{V}=D^{CVC}_{Y^{*}}$, and $X$ replaced by $X(\varepsilon)$, we get that the mapping $f_{\varepsilon}:X(\varepsilon)\rightarrow Y^{*}(\varepsilon)$ recovers the partitioning $\mathcal{V}$. That is, $D^{CVC}_{Y^{*}(\varepsilon)}$ is the same as $D^{CVC}_{Y^{*}}$. The proof is finished. $\square$
We demonstrate this by the following one-dimension example with two clusters, that is, $d=1$ and $K=2$. We uses the weighted matrix $W=(w_{ij})=E_{n}$ ( $E_{n_{1}}$ denotes the matrix of size $n_{1}\times n_{1}$ whose elements are all ones) and $p=2$. We only consider adding perturbation to a specific data. The solution of (CVC) is obtained by running the algorithm semismooth Newton-CG augmented Lagrangian method (Ssnal)  [^2] https://www.polyu.edu.hk/ama/profile/dfsun//Codes/Statistical-Optimization/ in [31].
Example 3.
Let X=[0, 2, 10, 14]1×4X=[0,\ 2,\ 10,\ $\in\mathbb{R}^{1\times 4}$, as shown in Figure 1. The solution of convex clustering model in (CVC) is $Y^{*}=[1,\ 1,\ 12,\ 12]$, with decision function $\mathcal{V}=D^{CVC}_{Y^{*}}=\left\{\{1,2\},\{3,4\}\right\}.$
Figure 1: X=[0, 2, 10, 14]X=[0,\ 2,\ 10,\ 14]
We perturb only on $x_{3}$. Let $X(\varepsilon)=[0,\ 2,\ 17,\ 14]$, that is, $\varepsilon=[0,\ 0,\ 7,\ 0]$. Easily see that $\mu_{12}^{(1)}(\varepsilon)=0,\ \mu_{34}^{(2)}(\varepsilon)=0$. Moreover, $n_{\alpha}w_{ij}(\varepsilon)-u_{ij}^{(\alpha)}(\varepsilon)=n_{\alpha}>0,\ \text{for all}\ i,j\in I_{\alpha},\ \alpha=1,2$, and $\gamma_{\min}^{\varepsilon}=\max\left\{\frac{2}{2},\frac{3}{2}\right\}=\frac{3}{2}<\gamma_{\max}^{\varepsilon}=\frac{14.5}{4}=\frac{29}{8}$. Therefore, conditions (C1) and (C2) hold and $D^{CVC}_{Y^{*}(\varepsilon)}=D^{CVC}_{Y^{*}}$. In fact, the solution of (CVC) gives $Y^{*}(\varepsilon)=[1,\ 1,\ 15.5,\ 15.5]$.
Following Example 3, we can similarly calculate that for any $\varepsilon_{3}\in\left(-\frac{6}{5},\frac{38}{3}\right)$, that is, $x_{3}(\varepsilon)\in\left(\frac{44}{5},\frac{68}{3}\right)$, conditions (C1) and (C2) both hold, implying that the clustering results will not be changed. This is indeed the truth since one can verify it by eyesight (See Figure 2).
Figure 2: $x_{3}(\varepsilon)\in\left(\frac{44}{5},\frac{68}{3}\right)$
In fact, if $\varepsilon$ does not satisfy (C1) or (C2), it is very likely that the clustering result will change compared with the unperturbed clustering result. Below we give another example to show this phenomenon.
Example 4.
[…]
Figure 3: $X(\varepsilon)=[0,\ 2,\ -4,\ 14]$
The above example gives rise to another interesting question: how to choose $\varepsilon$ in order to make the clustering result change? In fact, we have the following necessary condition for changing the clustering results.
Theorem 3.
If $D^{CVC}_{Y^{*}(\varepsilon)}\neq D^{CVC}_{Y^{*}}$, then either (C1) or (C2) fails for perturbed data $X(\varepsilon)$ under partition $D^{CVC}_{Y^{*}}$.
Proof. Assume for contradiction that $\varepsilon$ satisfies both (C1) and (C2) with perturbed data $X(\varepsilon)$ under partition $D^{CVC}_{Y^{*}}$. By Theorem 2, it holds that $D^{CVC}_{Y^{*}(\varepsilon)}=D^{CVC}_{Y^{*}}$, implying that the clustering results for $X(\varepsilon)$ will remain the same as $X$, which is a contradiction. Therefore, the proof is finished. $\square$
To conclude this section, we give a short summary. We discussed the role of $\varepsilon$ in the perturbation of convex clustering. In Theorem 2, we identified conditions on $\varepsilon$ under which the clustering result remains the same. We also provided a necessary condition on $\varepsilon$ in order that the clustering result can be changed.
4 Bilevel Model for Adversarial Learning
In this part, we will reformulate adversarial learning by bilevel optimization models. For general modeling and algorithmic background on bilevel optimization, see [42, 43].
As we mentioned in Section 3, the perturbation $\varepsilon$ on $X$ may led to the changed decision function, which means that the learning model is robust. On the other hand, the perturbation $\varepsilon$ on $X$ may lead to the change in decision result $D_{Y^{*}(\varepsilon)}$. It has two implications. Firstly, it means that such kind of noise is not neglectable. One needs to either do the denoising process to get rid of such noise or to improve the training process to make the learning model more robust. Secondly, from data attacking point of view, attack happens in such situation. In this case, a question arises: how could we choose the perturbation $\varepsilon$ such that some attacking criteria is maximized or minimized? This leads to the following two models for adversarial learning. The first model is to make the output of the perturbed learning model as huge difference as possible compared to the original learning model, which is described as ($a>0$ is given)
$\begin{split}\max_{\varepsilon\in\mathcal{X}}&\ U\left(\varepsilon\right)\\ \text{s.t. }&\ Y^{*}(\varepsilon)\in S(\varepsilon),\\ &\ \|\varepsilon\|\leq a,\end{split}$ (BL1)
where $U(\cdot):\mathcal{X}\to\mathbb{R}$ is defined to be the deviation function which measures the effect of attack, that is, the changes in decision $D_{Y^{*}(\varepsilon)}$ compared with the decision $D_{Y^{*}}$. To make $U(\cdot)$ well representing the effect of attack, $U(\cdot)$ must have the following properties:
  • (i)
    $U\left(\varepsilon\right)$ should be a nondecreasing function with respect to $\|\varepsilon\|$.
  • (ii)
    $U\left(\varepsilon\right)\geq 0$ for any $\varepsilon\in\mathcal{X}$, in particular, $U\left(\varepsilon\right)>0$ if $\varepsilon\neq 0$.
  • (iii)
    $U\left(\varepsilon\right)=0$. That is, if there is no perturbation ($\varepsilon$=0), the deviation should be zero.
In (BL1), $\|\varepsilon\|\leq a$ controls the magnitude of $\varepsilon$ such that it would not be too large, otherwise, the perturbation will be recognized, and the attack will fail. One can see that (BL1) is a bilevel optimization problem where $Y^{*}(\varepsilon)\in S(\varepsilon)$ describes the lower level problem, saying that $Y^{*}(\varepsilon)$ must be the solution of (Pε).
The second model is to minimize the scale of perturbation $\varepsilon$, such that the attacking effect reaches the prescribed effect level $\delta_{0}>0$. That is,
$\begin{split}\min_{\varepsilon\in\mathcal{X}}&\ \|\varepsilon\|\\ \text{s.t. }&\ Y^{*}(\varepsilon)\in S(\varepsilon),\\ &\ U\left(\varepsilon\right)\geq\delta_{0}.\end{split}$ (BL2)
Based on the two base bilevel models (BL1) and (BL2), one can consider some variants for adversarial learning. One typical example is to combine the deviation function and the scale of $\varepsilon$ by a penalty function, that is,
$\begin{split}\min_{\varepsilon\in\mathcal{X}}&\ U\left(\varepsilon\right)+\rho\|\varepsilon\|\\ \text{s.t. }&\ Y^{*}(\varepsilon)\in S(\varepsilon).\end{split}$
By choosing different weight $\rho>0$, one can balance the deviation function and the scale of the perturbation according to the user’s preference.
For example, the bilevel model for convex clustering is given as follows
$\begin{split}\min_{\varepsilon\in\mathcal{X}}&\ U\left(\varepsilon\right)+\rho\|\varepsilon\|\\ \text{s.t. }&\ Y^{*}(\varepsilon)=\arg\min_{Y\in\mathcal{Y}}L^{CVC}(X(\varepsilon),Y)\end{split}$
Since $L^{CVC}(\cdot,\cdot)$ is given, so it is white-box adversaial attack. How to solve this bilevel problem heavily depends on the choices of the deviation function $U(\cdot)$.
Here we would like to highlight the following two important issues.
  • (i)
    Firstly, due to different learning models, $U(\cdot)$ is usually difficult to design. As far as we know, there is little work focusing on developing efficient deviation function, which is in fact important to the two adversarial learning models. We will address this question in the scenario of clustering, which can be found in Section 5.
  • (ii)
    Secondly, due to the potential discontinuity of decision function $D$ and the possibly complicated learning models, it is usually difficult to solve the adversarial models (BL1) or (BL2), which is out of the scope of this paper and will be further discussed in our future work.
5 A Case Study on Deviation Functions
In this part, we will study the $\delta$-measure function in asversarial learning for clustering problems, to verify whether it is a deviation function.
As we mentioned above, due to the different structures of learning models as well as the variants of decision functions, the deviation function on the decision function can be in many different forms. Taking the binary classification as an example, the deviation function can be chosen as the norm of difference of the classification results, i.e.,
$U\left(\varepsilon\right)=\sum_{i=1}^{n}\|D^{SVM}_{Y^{*}(\varepsilon)}\left(x_{i}(\varepsilon)\right)-D^{SVM}_{Y^{*}}(x_{i})\|_{p}$
where $\|\cdot\|_{p}$ is $l_{p}$ norm ($1\leq p\leq\infty$) or $\|\cdot\|_{0}$, which counts the nonzero elements of a vector.
For clustering problems, recall the aim of clustering is to partition the data points into different groups, i.e., $D^{CVC}_{Y^{*}(\varepsilon)}$ is a partition of points in $X(\varepsilon)$, there are many different ways of measuring the clustering results. See [44, 45] for some of the clustering functions. One natural way is to use a matrix to represent the partition of points. Take $\mathcal{V}=\{V_{1},\cdots,V_{K}\}$ as an example. A 0-1 matrix $\widehat{D}(X)\in\mathbb{R}^{n\times K}$ is defined as follows:
$\widehat{D}(X)_{ij}=\begin{cases}1,\ \ &\ \text{if}\ x_{i}\in V_{j},\\ 0,\ \ &\ \text{otherwise}.\end{cases}$
Then the matrix $\widehat{D}(X)\widehat{D}(X)^{T}$ actually shows whether data points are grouped together, where
$\left(\widehat{D}(X)\widehat{D}(X)^{T}\right)_{ij}=\begin{cases}1,\ \ &\ \text{if}\ x_{i},x_{j}\ \text{are in the same partition},\\ 0,\ \ &\ \text{otherwise}.\end{cases}$
The following function is proposed by Biggio et al. [46]
$\delta(\varepsilon)=\left\|\widehat{D}(Y^{*}(\varepsilon))\widehat{D}(Y^{*}(\varepsilon))^{T}-\widehat{D}(Y^{*})\widehat{D}(Y^{*})^{T}\right\|_{F}^{2}.$ (2)
Chhabra et al. [5] believed that $\delta$ increases with the number of points that spill over from partition $V_{1}$ to $V_{2}$ for $K=2$. However, it is still not quite clear about whether this function can fully represent the deviation of $D_{Y^{*}(\varepsilon)}$ over the original $D_{Y^{*}}$. Therefore, below we conduct a systematic analysis on the property of $\delta$ defined in (2). We consider the following two scenarios.
5.1 Analysis on 2-Way Clustering
$K=2$, with $X\in\mathbb{R}^{d\times n}$ clustered into $V_{1}$ and $V_{2}$ with $|V_{1}|=n_{1},\ |V_{2}|=n_{2}$. Let $n=n_{1}+n_{2}$. Assume that under perturbation $\varepsilon$, the clustering of $X(\varepsilon)$ is changed to $V_{1}\setminus S$, $V_{2}\cup S$, where $S\subseteq V_{1}$, $|S|=s$. We have the following result.
Theorem 4.
For 2-way clustering, let $\delta\left(\varepsilon\right)$ be defined by (2).
  • (i)
    $\delta\left(\varepsilon\right)=2s(n-s)$.
  • (ii)
    If $s<\min(n_{1},\lceil\frac{n}{2}\rceil)$ [^3] Here $\lceil a\rceil$ denotes the smallest integer that is greater than or equal to $a$., $\delta\left(\varepsilon\right)$is a deviation function.
Proof: For simplicity, we use $\widehat{D}(\varepsilon)$ and $\widehat{D}(0)$ to represent $\widehat{D}(Y^{*}(\varepsilon))$ and $\widehat{D}(Y^{*})$. For (i), without loss of generality, let $V_{1}=\{x_{1},\cdots,x_{n_{1}}\},\ V_{2}=\{x_{n_{1}+1},\cdots,x_{n}\}$. Then it holds that
$\widehat{D}(0)=\begin{bmatrix}e_{n_{1}}&0\\ 0&e_{n_{2}}\end{bmatrix}\in\mathbb{R}^{n\times 2},\ \ \ \widehat{D}(0)\widehat{D}(0)^{T}=\begin{bmatrix}E_{n_{1}}&0\\ 0&E_{n_{2}}\end{bmatrix}\in\mathbb{R}^{n\times n},$
where $e_{n_{1}}$ is the column vector of length $n_{1}$ whose elements are all ones and $E_{n_{1}}$ denotes the matrix of size $n_{1}\times n_{1}$ whose elements are all ones. Here ’0’ denotes the zero vector or matrix of proper sizes. By changing the last $s$ data points in $V_{1}$ to $V_{2}$, we have
$\widehat{D}(\varepsilon)=\begin{bmatrix}e_{n_{1}-s}&0\\ 0&e_{n_{2}+s}\end{bmatrix}\in\mathbb{R}^{n\times 2},\ \ \ \widehat{D}(\varepsilon)\widehat{D}(\varepsilon)^{T}=\begin{bmatrix}E_{n_{1}-s}&0\\ 0&E_{n_{2}-s}\end{bmatrix}\in\mathbb{R}^{n\times n},$
leading to the following ($E_{i\times j}$ denotes the matrix of $i$ by $j$ with all elements one)
$\widehat{D}(0)\widehat{D}(0)^{T}-\widehat{D}(\varepsilon)\widehat{D}(\varepsilon)^{T}=\begin{bmatrix}0&E_{(n_{1}-s)\times s}&0\\ E_{s\times(n_{1}-s)}&0&-E_{s\times n_{2}}\\ 0&-E_{n_{2}\times s}&0\end{bmatrix}$
and
$\delta\left(\varepsilon\right)=\|\widehat{D}(0)\widehat{D}(0)^{T}-\widehat{D}(\varepsilon)\widehat{D}(\varepsilon)^{T}\|_{F}^{2}=2(n-s)s.$
This gives (i).
To show (ii), obviously, $\delta(0)=0$ and $\delta\left(\varepsilon\right)>0$ for any $Y^{*}(\varepsilon)\neq Y^{*}(0)$. Moreover, note that $0<s<n_{1}$, therefore, as $s$ increases for $s\in(0,\min(n_{1},\lceil\frac{n}{2}\rceil))$, $\delta(\cdot)$ is a nondecreasing function with respect to $s$. In other words, only when $s\in(0,\min(n_{1},\lceil\frac{n}{2}\rceil))$, $\delta$ is a deviation function with respect to $s$. $\square$
We give some examples as follows.
Example 5.
Let $X=[x_{1},\cdots,x_{5}]$ with partition $V_{1}=\{x_{1},\ x_{2},\ x_{3},\ x_{4}\}$ and $V_{2}=\{x_{5}\}$, which means $n_{1}=4$ and $n_{2}=1$. It holds that
$\widehat{D}(0)=\begin{bmatrix}e_{4}&0\\ 0&e_{1}\end{bmatrix}\in\mathbb{R}^{5\times 2}.$
Changing $x_{4}$ from $V_{1}$ to $V_{2}$, we get $V_{1}^{\prime}=\{x_{1},\ x_{2},\ x_{3}\}$ and $V_{2}^{\prime}=\{x_{4},\ x_{5}\}$, leading to the following
$\widehat{D}(\varepsilon)=\begin{bmatrix}e_{3}&0\\ 0&e_{2}\end{bmatrix}$.
Therefore, $\delta\left(\varepsilon\right)=\|\widehat{D}(0)\widehat{D}(0)^{T}-\widehat{D}(\varepsilon)\widehat{D}(\varepsilon)^{T}\|_{F}^{2}=8.$
Example 6.
Let $X$ and $V_{1},\ V_{2}$ be the same as in Example 5. Changing $x_{3},\ x_{4}$ from $V_{1}$ to $V_{2}$, to get $V_{1}^{\prime}=\{x_{1},\ x_{2}\}$ and $V_{2}^{\prime}=\{x_{3},\ x_{4},\ x_{5}\}$, we get
$\widehat{D}(\varepsilon)=\begin{bmatrix}e_{2}&0\\ 0&e_{3}\end{bmatrix}$,
which gives $\delta\left(\varepsilon\right)=12$.
Example 7.
Let $X$ and $V_{1},\ V_{2}$ be the same as in Example 5. Changing $x_{2},\ x_{3},\ x_{4}$ from $V_{1}$ to $V_{2}$ to get $V_{1}^{\prime}=\{x_{1}\}$ and $V_{2}^{\prime}=\{x_{2},\ x_{3},\ x_{4},\ x_{5}\}$, we get
$\widehat{D}(\varepsilon)=\begin{bmatrix}e_{1}&0\\ 0&e_{4}\end{bmatrix}$,
which also leads to $\delta\left(\varepsilon\right)=12$.
Comparing Example 6 and Example 7, the number of changed data is increasing. However, the deviation function $\delta$ is the same. In this case, $s=3>\lceil\frac{n}{2}\rceil$ in Example 7, implying that $\delta$ can not fully represent the deviation of the perturbed clustering results over the original clustering results.
In summary, Theorem 4 as well as the above examples show that choosing a proper deviation function $U$ is very important. However, sometimes it is difficult and even tricky to choose a good function which satisfy deviation properties (i)-(iii). Also, when dealing with adversarial learning models, we have to be very careful in order to choose a good deviation function since the chosen function may not fully reflect the changes in perturbation decision function after perturbation.
5.2 Analysis on 3-Way Clustering
$K=3$, with $X\in\mathbb{R}^{d\times n}$ clustered into $V_{1},\ V_{2},\ V_{3}$ with $|V_{i}|=n_{i},\ i=1,2,3$, and $n=\sum_{i=1}^{3}n_{i}$. After perturbation, the cluster changes to $V_{1}\setminus(S_{1}\cup S_{2})$, $V_{2}\cup S_{1}$, $V_{3}\cup S_{2}$, where $|S_{1}|=s_{1}$, $|S_{2}|=s_{2}$.
Theorem 5.
For 3-way clustering, it holds that
$\delta\left(\varepsilon\right)=(s_{1}+s_{2})\left(2n_{1}-(s_{1}+s_{2})\right)+s_{1}(2n_{2}-s_{1})+s_{2}(2n_{3}-s_{2}).$
In particular,
  • (i)
    If $S_{2}=\emptyset$, $\delta\left(\varepsilon\right)=2s_{1}(n-s_{1})$. Then $\delta(\cdot)$ is a deviation function.
  • (ii)
    If $n_{1}=n_{2}=n_{3}$ and $s_{1}=s_{2}$, $\delta\left(\varepsilon\right)=2s_{1}\left(\frac{4}{3}n-3s_{1}\right)$. Then $\delta(\cdot)$ is a deviation function.
  • (iii)
    If $s_{1}=s_{2}$, $\delta\left(\varepsilon\right)=s_{1}(4n_{1}+2n_{2}+2n_{3}-6s_{1})$. For $s_{1}\in\left(0,\min\left(\lceil\frac{2n_{1}+n_{2}+n_{3}}{6}\rceil,\lceil\frac{n_{1}}{2}\rceil\right)\right)$, $\delta(\cdot)$ is a deviation function.
Proof. Without loss of generality, assume that $V_{1}=\{x_{1},\cdots,x_{n_{1}}\}$, $V_{2}=\{x_{n_{1}+1},\cdots,x_{n_{1}+n_{2}}\}$, $V_{3}=\{x_{n_{1}+n_{2}+1},\cdots,x_{n}\}$. It holds that
$\widehat{D}(0)=\begin{bmatrix}e_{n_{1}}&0&0\\ 0&e_{n_{2}}&0\\ 0&0&e_{n_{3}}\end{bmatrix}\in\mathbb{R}^{n\times 3},\ \ \ \widehat{D}(0)\widehat{D}(0)^{T}=\begin{bmatrix}E_{n_{1}}&0&0\\ 0&E_{n_{2}}&0\\ 0&0&E_{n_{3}}\end{bmatrix}\in\mathbb{R}^{n\times n}.$
Now a subset of data points $S_{1}\subseteq V_{1}$ changes their cluster membership from $V_{1}$ to $V_{2}$ and a subset of data points $S_{2}\subseteq V_{1}$ changes their cluster membership from $V_{1}$ to $V_{3}$. It holds that
$\widehat{D}(\varepsilon)=\begin{bmatrix}e_{n_{1}-(s_{1}+s_{2})}&0&0\\ 0&e_{s_{1}}&0\\ 0&0&e_{s_{2}}\\ 0&e_{n_{2}}&0\\ 0&0&e_{n_{3}}\\ \end{bmatrix}\in\mathbb{R}^{n\times 3},\ \ \widehat{D}(\varepsilon)\widehat{D}(\varepsilon)^{T}=\begin{bmatrix}E_{n_{1}-(s_{1}+s_{2})}&0&0&0&0\\ 0&E_{s_{1}}&0&E_{s_{1}\times n_{2}}&0\\ 0&0&E_{s_{2}}&0&E_{s_{2}\times n_{3}}\\ 0&E_{n_{2}\times s_{1}}&0&E_{n_{2}}&0\\ 0&0&E_{n_{3}\times s_{2}}&0&E_{n_{3}}\\ \end{bmatrix}\in\mathbb{R}^{n\times n},$
leading to the following
$\hat{D}(0)\hat{D}(0)^{T}-\hat{D}(\varepsilon)\hat{D}(\varepsilon)^{T}=\begin{bmatrix}0&E_{(n_{1}-(s_{1}+s_{2}))\times s_{1}}&E_{(n_{1}-(s_{1}+s_{2}))\times s_{2}}&0&0\\ E_{s_{1}\times(n_{1}-(s_{1}+s_{2}))}&0&E_{s_{1}\times s_{2}}&-E_{s_{1}\times n_{2}}&0\\ E_{s_{2}\times(n_{1}-(s_{1}+s_{2}))}&E_{s_{2}\times s_{1}}&0&0&-E_{s_{2}\times n_{3}}\\ 0&-E_{n_{2}\times s_{1}}&0&0&0\\ 0&0&-E_{n_{3}\times s_{2}}&0&0\\ \end{bmatrix}$.
After calculation, we get
$\begin{split}\delta(\varepsilon)&=\|\widehat{D}(0)\widehat{D}(0)^{T}-\widehat{D}(\varepsilon)\widehat{D}(\varepsilon)^{T}\|_{F}^{2}\\ &=(n_{1}-(s_{1}+s_{2}))(s_{1}+s_{2})+s_{1}(n_{1}-s_{1}+n_{2})+s_{2}(n_{1}-s_{2}+n_{3})+n_{2}s_{1}+n_{3}s_{2}\\ &=2(n_{1}(s_{1}+s_{2})+n_{2}s_{1}+n_{3}s_{2})-((s_{1}+s_{2})^{2}+s_{1}^{2}+s_{2}^{2})\\ &=(s_{1}+s_{2})(2n_{1}-(s_{1}+s_{2}))+s_{1}(2n_{2}-s_{1})+s_{2}(2n_{3}-s_{2}).\end{split}$
This gives the first part of the results.
For (i), $S_{2}=\emptyset$. In this case, the third cluster $V_{3}$ does not play any role. Then $\delta(\cdot)$ reduces to $2s_{1}(n_{1}+n_{2}-s_{1})$, which coincides with the results in Theorem 4.
For (ii), if $n_{1}=n_{2}=n_{3}$, $s_{1}=s_{2}$, $\delta(\cdot)$ takes the following form (note that $s=s_{1}+s_{2}=2s_{1}$)
$\delta(\varepsilon)=2s_{1}(4n_{1}-3s_{1})=2s_{1}\left(\frac{4}{3}n-3s_{1}\right).$
Note that for $s_{1}\in\left(0,\lceil\frac{2n}{9}\rceil\right),\ \delta$ is nondecreasing. Also note that $s_{1}<\lceil\frac{n_{1}}{2}\rceil=\lceil\frac{n}{6}\rceil$, therefore, for any $s_{1}$ in that case, $s_{1}<\lceil\frac{2n}{9}\rceil$, which means that $\delta(\cdot)$ is always nondecreasing. That is, for any $s_{1}$, $\delta(\cdot)$ is always a deviation function.
For (iii), if $s_{1}=s_{2}$, then
[…]
So for $s_{1}\in\left(0,\min\left(\lceil\frac{2n_{1}+n_{2}+n_{3}}{6}\rceil,\lceil\frac{n_{1}}{2}\rceil\right)\right),\ \delta(\cdot)$ is nondecreasing. Therefore, it is a deviation function. The proof is finished. $\square$
Below we show some examples (Example 8 and 9) for case (ii) and case (iii) (Example 10-12).
Example 8.
Let $X=[x_{1},\cdots,x_{15}]$ with partition $V_{1}=\{x_{1},\cdots,x_{5}\}$, $V_{2}=\{x_{6},\cdots,x_{10}\}$ and $V_{3}=\{x_{11},\cdots,x_{15}\}$, that is, $n_{1}=n_{2}=n_{3}=5$. It holds that
$\widehat{D}(0)=\begin{bmatrix}e_{5}&0&0\\ 0&e_{5}&0\\ 0&0&e_{5}\end{bmatrix}\in\mathbb{R}^{15\times 3},\ \ \ \widehat{D}(0)\widehat{D}(0)^{T}=\begin{bmatrix}E_{5}&0&0\\ 0&E_{5}&0\\ 0&0&E_{5}\end{bmatrix}$.
Changing $x_{2}$ from $V_{1}$ to $V_{2}$ and $x_{3}$ from $V_{1}$ to $V_{3}$, which means $s_{1}=s_{2}=1$, we get
$\widehat{D}(\varepsilon)=\begin{bmatrix}e_{3}&0&0\\ 0&e_{1}&0\\ 0&0&e_{1}\\ 0&e_{5}&0\\ 0&0&e_{5}\\ \end{bmatrix},\ \ \ \widehat{D}(\varepsilon)\widehat{D}(\varepsilon)^{T}=\begin{bmatrix}E_{3}&0&0&0&0\\ 0&1&0&E_{1\times 5}&0\\ 0&0&1&0&E_{1\times 5}\\ 0&E_{5\times 1}&0&E_{5}&0\\ 0&0&E_{5\times 1}&0&E_{5}\\ \end{bmatrix}$.
It holds that
$\widehat{D}(0)\widehat{D}(0)^{T}-\widehat{D}(\varepsilon)\widehat{D}(\varepsilon)^{T}=\begin{bmatrix}0&E_{3\times 1}&E_{3\times 1}&0&0\\ E_{1\times 3}&0&1&-E_{1\times 5}&0\\ E_{1\times 3}&1&0&0&-E_{1\times 5}\\ 0&-E_{5\times 1}&0&0&0\\ 0&0&-E_{5\times 1}&0&0\\ \end{bmatrix}$,
and $\delta\left(\varepsilon\right)=\|\widehat{D}(0)\widehat{D}(0)^{T}-\widehat{D}(\varepsilon)\widehat{D}(\varepsilon)^{T}\|_{F}^{2}=34.$
Example 9.
Let $X$ and $V_{1},\ V_{2},\ V_{3}$ be the same as in Example 8. $\widehat{D}(0)$ is the same as in Example 8. Changing $x_{2},\ x_{3}$ from $V_{1}$ to $V_{2}$ and $x_{4},\ x_{5}$ from $V_{1}$ to $V_{3}$, which means $s_{1}=s_{2}=2$, we get
$\widehat{D}(\varepsilon)=\begin{bmatrix}e_{1}&0&0\\ 0&e_{2}&0\\ 0&0&e_{2}\\ 0&e_{5}&0\\ 0&0&e_{5}\\ \end{bmatrix},\ \ \ \widehat{D}(\varepsilon)\widehat{D}(\varepsilon)^{T}=\begin{bmatrix}1&0&0&0&0\\ 0&E_{2}&0&E_{2\times 5}&0\\ 0&0&E_{2}&0&E_{2\times 5}\\ 0&E_{5\times 2}&0&E_{5}&0\\ 0&0&E_{5\times 2}&0&E_{5}\\ \end{bmatrix}$.
It holds that
$\widehat{D}(0)\widehat{D}(0)^{T}-\widehat{D}(\varepsilon)\widehat{D}(\varepsilon)^{T}=\begin{bmatrix}0&E_{1\times 2}&E_{1\times 2}&0&0\\ E_{2\times 1}&0&E_{2}&-E_{2\times 5}&0\\ E_{2\times 1}&E_{2}&0&0&-E_{2\times 5}\\ 0&-E_{5\times 2}&0&0&0\\ 0&0&-E_{5\times 2}&0&0\\ \end{bmatrix}$,
and $\delta\left(\varepsilon\right)=\|\widehat{D}(0)\widehat{D}(0)^{T}-\widehat{D}(\varepsilon)\widehat{D}(\varepsilon)^{T}\|_{F}^{2}=56.$
For Example 8 and Example 9, it can be noticed that for case (ii), if $s_{1}$ increases, then $\delta(\cdot)$ increases, which fully reflect the changes of $\varepsilon$ in $\delta$-measure function.
Example 10.
Let $X=[x_{1},\cdots,x_{11}]$ with partition $V_{1}=\{x_{1},\cdots,x_{9}\}$, $V_{2}=\{x_{10}\}$ and $V_{3}=\{x_{11}\}$, i.e., $n_{1}=9$, $n_{2}=1$ and $n_{3}=1$. It holds that
$\widehat{D}(0)=\begin{bmatrix}e_{9}&0&0\\ 0&e_{1}&0\\ 0&0&e_{1}\end{bmatrix}\in\mathbb{R}^{11\times 3},\ \ \ \widehat{D}(0)\widehat{D}(0)^{T}=\begin{bmatrix}E_{9}&0&0\\ 0&1&0\\ 0&0&1\end{bmatrix}$.
Changing $x_{6},\ x_{7}$ from $V_{1}$ to $V_{2}$ and $x_{8},\ x_{9}$ from $V_{1}$ to $V_{3}$, which means $s_{1}=s_{2}=2$, we get
$\widehat{D}(\varepsilon)=\begin{bmatrix}e_{5}&0&0\\ 0&e_{2}&0\\ 0&0&e_{2}\\ 0&e_{1}&0\\ 0&0&e_{1}\end{bmatrix},\ \ \ \widehat{D}(\varepsilon)\widehat{D}(\varepsilon)^{T}=\begin{bmatrix}E_{5}&0&0&0&0\\ 0&E_{2}&0&E_{2\times 1}&0\\ 0&0&E_{2}&0&E_{2\times 1}\\ 0&E_{1\times 2}&0&E_{1}&0\\ 0&0&E_{1\times 2}&0&E_{1}\end{bmatrix}$.
It holds that
$\widehat{D}(0)\widehat{D}(0)^{T}-\widehat{D}(\varepsilon)\widehat{D}(\varepsilon)^{T}=\begin{bmatrix}0&E_{5\times 2}&E_{5\times 2}&0&0\\ E_{2\times 5}&0&E_{2}&-E_{2\times 1}&0\\ E_{2\times 5}&E_{2}&0&0&-E_{2\times 1}\\ 0&-E_{1\times 2}&0&0&0\\ 0&0&-E_{1\times 2}&0&0\end{bmatrix}$,
and
$\delta\left(\varepsilon\right)=\|\widehat{D}(0)\widehat{D}(0)^{T}-\widehat{D}(\varepsilon)\widehat{D}(\varepsilon)^{T}\|_{F}^{2}=56.$
Example 11.
Let $X$ and $V_{1},\ V_{2},\ V_{3}$ be the same as in Example 10. $\widehat{D}(0)$ is the same as in Example 10. Changing $x_{4},\ x_{5},\ x_{6}$ from $V_{1}$ to $V_{2}$ and $x_{7},\ x_{8},\ x_{9}$ from $V_{1}$ to $V_{3}$, which means $s_{1}=s_{2}=3$, we get
$\widehat{D}(\varepsilon)=\begin{bmatrix}e_{3}&0&0\\ 0&e_{3}&0\\ 0&0&e_{3}\\ 0&e_{1}&0\\ 0&0&e_{1}\end{bmatrix},\qquad\widehat{D}(\varepsilon)\widehat{D}(\varepsilon)^{T}=\begin{bmatrix}E_{3}&0&0&0&0\\ 0&E_{3}&0&E_{3\times 1}&0\\ 0&0&E_{3}&0&E_{3\times 1}\\ 0&E_{1\times 3}&0&E_{1}&0\\ 0&0&E_{1\times 3}&0&E_{1}\end{bmatrix}$.
It holds that
$\widehat{D}(0)\widehat{D}(0)^{T}-\widehat{D}(\varepsilon)\widehat{D}(\varepsilon)^{T}=\begin{bmatrix}0&E_{3\times 3}&E_{3\times 3}&0&0\\ E_{3\times 3}&0&E_{3}&-E_{3\times 1}&0\\ E_{3\times 3}&E_{3}&0&0&-E_{3\times 1}\\ 0&-E_{1\times 3}&0&0&0\\ 0&0&-E_{1\times 3}&0&0\end{bmatrix}$,
and
$\delta\left(\varepsilon\right)=\|\widehat{D}(0)\widehat{D}(0)^{T}-\widehat{D}(\varepsilon)\widehat{D}(\varepsilon)^{T}\|_{F}^{2}=66.$
Example 12.
Let $X$ and $V_{1},\ V_{2},\ V_{3}$ be the same as in Example 10. $\widehat{D}(0)$ is the same as in Example 10. Changing $\{x_{2},\cdots,x_{5}\}$ from $V_{1}$ to $V_{2}$ and $\{x_{6},\cdots,x_{9}\}$ from $V_{1}$ to $V_{3}$, which means $s_{1}=s_{2}=4$, we get
$\widehat{D}(\varepsilon)=\begin{bmatrix}e_{1}&0&0\\ 0&e_{4}&0\\ 0&0&e_{4}\\ 0&e_{1}&0\\ 0&0&e_{1}\end{bmatrix},\ \ \ \widehat{D}(\varepsilon)\widehat{D}(\varepsilon)^{T}=\begin{bmatrix}E_{1}&0&0&0&0\\ 0&E_{4}&0&E_{4\times 1}&0\\ 0&0&E_{4}&0&E_{4\times 1}\\ 0&E_{1\times 4}&0&E_{1}&0\\ 0&0&E_{1\times 4}&0&E_{1}\end{bmatrix}$.
It holds that
$\widehat{D}(0)\widehat{D}(0)^{T}-\widehat{D}(\varepsilon)\widehat{D}(\varepsilon)^{T}=\begin{bmatrix}0&E_{1\times 4}&E_{1\times 4}&0&0\\ E_{4\times 1}&0&E_{4}&-E_{4\times 1}&0\\ E_{4\times 1}&E_{4}&0&0&-E_{4\times 1}\\ 0&-E_{1\times 4}&0&0&0\\ 0&0&-E_{1\times 4}&0&0\end{bmatrix}$,
and
$\delta\left(\varepsilon\right)=\|\widehat{D}(0)\widehat{D}(0)^{T}-\widehat{D}(\varepsilon)\widehat{D}(\varepsilon)^{T}\|_{F}^{2}=64.$
Comparing Example 10 and Example 11, $s_{1}<\min\left(\lceil\frac{2n_{1}+n_{2}+n_{3}}{6}\rceil,\lceil\frac{n_{1}}{2}\rceil\right)=4$, so it increases as the number of changed data points grows. However, when comparing Example 11 and Example 12, the deviation function $\delta(\cdot)$ decreases. In this case, $s_{1}=4\notin(0,4)$ in Example 12, implying that $\delta(\cdot)$ cannot fully represent the deviation of the perturbed clustering results from the original clustering results.
However, if $n_{1},\ n_{2},\ n_{3}$ are not the same or $s_{1}\neq s_{2}$, the situation is more complicated to analyze. Some small examples are given below, as shown in Example 13, Example 14 and Example 15.
Example 13.
Let $X$ and $V_{1},\ V_{2},\ V_{3}$ be the same as in Example 10. $\widehat{D}(0)$ is the same as in Example 10. Changing $x_{4},\ x_{5}$ from $V_{1}$ to $V_{2}$ and $\{x_{6},\cdots,x_{9}\}$ from $V_{1}$ to $V_{3}$, which means $s_{1}=2,\ s_{2}=4$, we get
$\widehat{D}(\varepsilon)=\begin{bmatrix}e_{3}&0&0\\ 0&e_{2}&0\\ 0&0&e_{4}\\ 0&e_{1}&0\\ 0&0&e_{1}\end{bmatrix},\ \ \ \widehat{D}(\varepsilon)\widehat{D}(\varepsilon)^{T}=\begin{bmatrix}E_{3}&0&0&0&0\\ 0&E_{2}&0&E_{2\times 1}&0\\ 0&0&E_{4}&0&E_{4\times 1}\\ 0&E_{1\times 2}&0&E_{1}&0\\ 0&0&E_{1\times 4}&0&E_{1}\end{bmatrix}$.
It holds that
$\widehat{D}(0)\widehat{D}(0)^{T}-\widehat{D}(\varepsilon)\widehat{D}(\varepsilon)^{T}=\begin{bmatrix}0&E_{3\times 2}&E_{3\times 4}&0&0\\ E_{2\times 3}&0&E_{2}&-E_{2\times 1}&0\\ E_{4\times 3}&E_{4}&0&0&-E_{4\times 1}\\ 0&-E_{1\times 2}&0&0&0\\ 0&0&-E_{1\times 4}&0&0\end{bmatrix}$,
and $\delta\left(\varepsilon\right)=64.$
Example 14.
Let $X$ and $V_{1},\ V_{2},\ V_{3}$ be the same as in Example 10. $\widehat{D}(0)$ is the same as in Example 10. Changing $x_{3},\ x_{4}$ from $V_{1}$ to $V_{2}$ and $\{x_{5},\cdots,x_{9}\}$ from $V_{1}$ to $V_{3}$, which means $s_{1}=2,\ s_{2}=5$, we get
$\widehat{D}(\varepsilon)=\begin{bmatrix}e_{2}&0&0\\ 0&e_{2}&0\\ 0&0&e_{5}\\ 0&e_{1}&0\\ 0&0&e_{1}\end{bmatrix},\ \ \ \widehat{D}(\varepsilon)\widehat{D}(\varepsilon)^{T}=\begin{bmatrix}E_{2}&0&0&0&0\\ 0&E_{2}&0&E_{2\times 1}&0\\ 0&0&E_{5}&0&E_{5\times 1}\\ 0&E_{1\times 2}&0&E_{1}&0\\ 0&0&E_{1\times 5}&0&E_{1}\end{bmatrix}$.
It holds that
$\widehat{D}(0)\widehat{D}(0)^{T}-\widehat{D}(\varepsilon)\widehat{D}(\varepsilon)^{T}=\begin{bmatrix}0&E_{2\times 2}&E_{2\times 5}&0&0\\ E_{2\times 2}&0&E_{2}&-E_{2\times 1}&0\\ E_{5\times 2}&E_{5}&0&0&-E_{5\times 1}\\ 0&-E_{1\times 2}&0&0&0\\ 0&0&-E_{1\times 5}&0&0\end{bmatrix}$,
and $\delta\left(\varepsilon\right)=62.$
Example 15.
Let $X$ and $V_{1},\ V_{2},\ V_{3}$ be the same as in Example 10. $\widehat{D}(0)$ is the same as in Example 10. Changing $x_{3},\ x_{4},\ x_{5}$ from $V_{1}$ to $V_{2}$ and $\{x_{6},\cdots,x_{9}\}$ from $V_{1}$ to $V_{3}$, which means $s_{1}=3,\ s_{2}=4$, we get
$\widehat{D}(\varepsilon)=\begin{bmatrix}e_{2}&0&0\\ 0&e_{3}&0\\ 0&0&e_{4}\\ 0&e_{1}&0\\ 0&0&e_{1}\end{bmatrix},\qquad\widehat{D}(\varepsilon)\widehat{D}(\varepsilon)^{T}=\begin{bmatrix}E_{2}&0&0&0&0\\ 0&E_{3}&0&E_{3\times 1}&0\\ 0&0&E_{4}&0&E_{4\times 1}\\ 0&E_{1\times 3}&0&E_{1}&0\\ 0&0&E_{1\times 4}&0&E_{1}\end{bmatrix}$.
It holds that
$\widehat{D}(0)\widehat{D}(0)^{T}-\widehat{D}(\varepsilon)\widehat{D}(\varepsilon)^{T}=\begin{bmatrix}0&E_{2\times 3}&E_{2\times 4}&0&0\\ E_{3\times 2}&0&E_{3}&-E_{3\times 1}&0\\ E_{4\times 2}&E_{4}&0&0&-E_{4\times 1}\\ 0&-E_{1\times 3}&0&0&0\\ 0&0&-E_{1\times 4}&0&0\end{bmatrix}$,
and $\delta\left(\varepsilon\right)=66.$
Comparing Example 13 with Example 14, we see that although more points are changed, $\delta(\cdot)$ decreases. It indicates that in this situation, $\delta$-measure function are not fully reflect the changes in partitioning. Moreover, comparing Example 14 with Example 15 shows that even with the same number of changed points ($s_{1}+s_{2}=7$), $\delta(\cdot)$ can differ because the points are reassigned to different clusters. This highlights that $\delta(\cdot)$ is influenced not only by how many points are perturbed, but also by how those perturbations are distributed across clusters.
Theorem 5 shows that the $\delta(\cdot)$ under 3-way depends on two factors: how many points leave $V_{1}$ in total and how unevenly they split between $V_{2}$ and $V_{3}$. In the balanced, symmetric case (Example 8 and Example 9), $\delta(\cdot)$ is a deviation function. When cluster sizes or perturbed sizes are unbalanced, the effect is more complicated. In summary, $\delta(\cdot)$ is affected by both the number of perturbed points and the asymmetry of the reassignment, and it is predictable under balanced, symmetric case.
6 Conclusions
In this paper, we proposed the unified bilevel models for adversarial learning. We investigated the adversarial attack in clustering models and interpreted it from data perturbation point of view. Taking clustering as an example, we interpreted the attack as the perturbation of data, and provided sufficient conditions on the robustness of the convex clustering model. Finally, we study the properties of deviation function and provide a concrete choice for the deviation function. However, there are still a lot of questions that need to be further investigated, such as how to solve the two bilevel models for adversarial learning. We leave them as future research topics.
Conflict of interest
The authors declare that they have no conflict of interest.